# On the Alexander polynomial of special alternating links

### v2

@@ 6.2. Definition of our weight-preserving bijection
  3. Case 3: The final vertex $\operatorname{fin}(f)=v_{e}$ for some edge $e\in\operatorname{E}(D)$ which is not in $T$, and the unique edge $f^{\prime}$ in the boundary of $C_{f}$ with $\operatorname{fin}(f)=\operatorname{init}(f^{\prime})$ appears before $f$ in the Eulerian tour $\Phi(T)$. We argue that this case would violate condition (a).
    Assuming Case 3, note that there must be a path $P$ in $\Phi(T)$ from $f^{\prime}$ to $f$, as shown below.
@@
    Suppose that $g\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ is an edge not in $P$ such that $\operatorname{fin}(g)\in\operatorname{V}(P)$. Let $g^{\prime}$ be the edge immediately following $g$ in $\Phi(T)$; in particular, $\operatorname{init}(g^{\prime})=\operatorname{fin}(g)$. By construction of $\Phi$, $g$ and $g^{\prime}$ must either both lie inside the region bounded by $P$ or both lie outside it (see Definition 6.3). This implies that once $\Phi(T)$ reaches $f$, it cannot re-enter the portion of the graph enclosed by $P$. In particular, no edge in the boundary of $C_{f}$ can be reached after $f$ in $\Phi(T)$, so we have a contradiction with condition (a).
Proof of Theorem 6.5. Let $w$ denote the degree of ${\rm wt}_{C}(F(T))$. For a region $R\in R(\overrightarrow{\mathcal{G}(L)})$, let $\partial R$ denote the oriented boundary of $R$ as a subgraph of $\overrightarrow{\mathcal{G}(L)}$.
@@
$=(|V(\overrightarrow{\mathcal{G}(L)})|-1)-\left|\{f\in\operatorname{E}(F(T))\ |\ f\ \text{is}\ \text{blue}\}\right|.$
Apply Lemma 6.8 to replace the second term in the difference above. Note that there are no edges of $\overrightarrow{\mathcal{G}(L)}$ which violate conditions $(a)$ and $(b)$ of the lemma simultaneously. Accordingly,
$w$ $=(|V(\overrightarrow{\mathcal{G}(L)})|-1)-\sum_{R\in R(\overrightarrow{\mathcal{G}(L)})\ \text{blue}}((|\operatorname{E}(\partial R)|-1)-\left|\{e\in\operatorname{E}(T)\ |\ v_{e}\in\partial R\ \text{and}\ e\ \text{points towards}\ r\}\right|)$
$=(|V(\overrightarrow{\mathcal{G}(L)})|-1)-\sum_{R\in R(\overrightarrow{\mathcal{G}(L)})\ \text{blue}}(|\operatorname{E}(\partial R)|-1)$
[…]
$=(|V(\overrightarrow{\mathcal{G}(L)})|-1)-\left|\{e\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})\ |\ e\ \text{is blue}\}\right|+m_{2}+\left|\{e\in\operatorname{E}(T)\ |\ e\ \text{points towards}\ r\}\right|.$
Now, the map $e\mapsto\operatorname{init}(e)$ is a bijection between blue edges of $\overrightarrow{\mathcal{G}(L)}$ and $\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$, so
$w=m_{2}-1+\left|\{e\in\operatorname{E}(F(T))\ |\ e\ \text{points towards}\ r\}\right|=m_{2}-1+|E(T)\backslash E_{\text{away}}(T)|.$
Therefore, ${\rm wt}_{C}(A)={\rm wt}_{K}(T)t^{m_{2}-1}$. ∎

### v3

@@ 6.2. Definition of our weight-preserving bijection
  2. Case 2: The final vertex $\operatorname{fin}(f)=v_{e}$ for some edge $e\in\operatorname{E}(D)$ which is not in $T$, and the unique edge $f^{\prime}$ in the boundary of $R$ with $\operatorname{fin}(f)=\operatorname{init}(f^{\prime})$ appears before $f$ in the Eulerian tour $\Phi(T)$.
    Assuming Case 2, note that there must be a path $P$ in $\Phi(T)$ from $f^{\prime}$ to $f$, as shown below.
@@
    Suppose that $g\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ is an edge not in $P$ such that $\operatorname{fin}(g)\in\operatorname{V}(P)$. Let $g^{\prime}$ be the edge immediately following $g$ in $\Phi(T)$; in particular, $\operatorname{init}(g^{\prime})=\operatorname{fin}(g)$. By construction of $\Phi$, $g$ and $g^{\prime}$ must either both lie inside the region bounded by $P$ or both lie outside it (see Definition 6.4). This implies that once $\Phi(T)$ reaches $f$, it cannot re-enter the portion of the graph enclosed by $P$. In particular, no edge in the boundary of $R$ can be reached after $f$ in $\Phi(T)$. This contradicts the assumption that $f$ is not the last edge bounding $R$ that is reached by $\Phi(T)$.
Lemma 6.12.
Fix a spanning tree $T$ of $D$, and let $r\in\operatorname{V}(D)$ denote the vertex corresponding to the exterior region of $L$. Fix $e_{0}\in\operatorname{E}(D)$ with $\operatorname{fin}(e_{0})=r$. Let $v=v_{e_{0}}\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$, and let $f_{v}\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ be the unique edge in the boundary of the exterior region such that $\operatorname{fin}(f_{v})=v$.
Let $f\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ be in the boundary of a blue region $R$, and suppose that $f$ is the last edge of $R$ to be reached by $\Phi(T)$. Then $\operatorname{fin}(f)\neq v_{e}$ for any $e\in T$.
Proof.
Suppose $\operatorname{fin}(f)=v_{e}$ for some $e\in T$. Let $f^{\prime}$ be the unique edge on the boundary of $R$ such that $\operatorname{init}(f^{\prime})=v_{e}$. Note that $f^{\prime}\neq f_{v}$. Otherwise, we would have $f^{\prime}=f_{e_{0}}$ where $\operatorname{fin}(e_{0})=r$. Lemma 6.10 would then imply that $f^{\prime}$ is gold, yielding a contradiction.
Since $f^{\prime}\neq f_{v}$, Definition 6.4 requires that $f^{\prime}$ immediately follows $f$ in $\Phi(T)$. This contradicts our assumption that $f$ is the last edge of $R$ to be reached by $\Phi(T)$. ∎
Proof of Theorem 6.7. Let $w$ denote the degree of ${\rm wt}_{C}(F(T))$. For a region $R\in R(\overrightarrow{\mathcal{G}(L)})$, let $\partial R$ denote the oriented boundary of $R$ as a subgraph of $\overrightarrow{\mathcal{G}(L)}$. Let $\mathcal{B}$ denote the set of blue edges of $\overrightarrow{\mathcal{G}(L)}$.
@@
$=(|\operatorname{V}(\overrightarrow{\mathcal{G}(L)})|-1)-\left|\{f\in\operatorname{E}(F(T))\ |\ f\ \text{is}\ \text{blue}\}\right|$
$=(|\operatorname{V}(\overrightarrow{\mathcal{G}(L)})|-1)-\left|\{f\in\operatorname{E}(F(T))\ |\ f=f_{e}\ \text{for some}\ e\in T\ \text{and}\ f\ \text{is}\ \text{blue}\}\right|$
|$\{f\in\operatorname{E}(F(T))\ |\ f\neq f_{e}\ \text{for any}\ e\in T\ \text{and}\ f\ \text{is}\ \text{blue}\}\right|$ |
We consider the four possible cases for each blue edge $f\in\operatorname{E}(\overrightarrow{\mathcal{G}}(L))$.
  1. Case 1: The final vertex $\operatorname{fin}(f)=v_{e}$ for some $e\in\operatorname{E}_{\text{away}}(T)$. In this case, we must have $f=f_{e}$. Otherwise, $f_{e}$ would be a gold edge (since there are only two edges of $\overrightarrow{\mathcal{G}(L)}$ with a given final vertex – one blue and one gold), and Lemma 6.10 would imply that $e$ is directed towards $r$, yielding a contradiction.
  2. Case 2: The final vertex $\operatorname{fin}(f)=v_{e}$ for some $e\in\operatorname{E}(T)\setminus\operatorname{E}_{\text{away}}(T)$. Then, $f$ cannot be an edge of $F(T)$, since by Lemma 6.10, $f_{e}$ must be gold.
  3. Case 3: The final vertex $\operatorname{fin}(f)\neq v_{e}$ for any $e\in T$, and $f$ is the last edge of the blue region it bounds reached by $\Phi(T)$. Then, by Lemma 6.11, $f$ cannot be an edge of $F(T)$.
  4. Case 4: The final vertex $\operatorname{fin}(f)\neq v_{e}$ for any $e\in T$, and $f$ is not the last edge of the blue region it bounds reached by $\Phi(T)$. In this case, Lemma 6.11 implies that $f\in\operatorname{E}(F(T))$.
The term |$\left|\{f\in\operatorname{E}(F(T))\ |\ f=f_{e}\ \text{for some}\ e\in T\ \text{and}\ f\ \text{is}\ \text{blue}\}\right|$ in our computation of $w$ precisely counts the blue edges falling into Case 1, so we may replace this term with $|\operatorname{E}_{\text{away}}(T)|$. Furthermore, the term |$\left|\{f\in\operatorname{E}(F(T))\ |\ f\neq f_{e}\ \text{for any}\ e\in T\ \text{and}\ f\ \text{is}\ \text{blue}\}\right|$ counts the blue edges falling into Case 4. By Lemma 6.12, we can replace this term with $|\mathcal{B}|-|\operatorname{E}(T)|-m_{2}$. We then have
$w$ $=(|\operatorname{V}(\overrightarrow{\mathcal{G}(L)})|-1)-|\mathcal{B}|+|\operatorname{E}(T)|-|\operatorname{E}_{\text{away}}(T)|+m_{2}$
$=(|\operatorname{V}(\overrightarrow{\mathcal{G}(L)})|-1)-|\mathcal{B}|+|\operatorname{E}(T)\setminus\operatorname{E}_{\text{away}}(T)|+m_{2}.$
Finally, the map $f\mapsto\operatorname{fin}(f)$ is a bijection from $\mathcal{B}$ to $\operatorname{V}(\overrightarrow{G}(L))$. Therefore,
$w=|\operatorname{E}(T)\setminus\operatorname{E}_{\text{away}}(T)|+m_{2}-1,$
and we can conclude that ${\rm wt}_{C}(A)={\rm wt}_{K}(T)t^{m_{2}-1}$.